OpenStudy (anonymous):

Expand (a+b) do the 4th power

jhonyy9 (jhonyy9):

(a+b)*4 = ((a+b)*2)*2 =(a*2 +2ab +b*2)*2 =(a*2 +2ab +b*2)(a*2 +2ab +b*2)

OpenStudy (nikvist):

\[(a+b)^4=a^4+4a^3b+6a^2b^2+4ab^3+b^4\]

OpenStudy (nikvist):

because \[(a+b)^n=\sum_{k=0}^n comb(n,k)\,\,a^{n-k}b^k\]